OpenStudy (anonymous):

how would you rite 19 over 25 as an percentage? thank you.

OpenStudy (anonymous):

\[\frac{ 19 }{ 25 }=\frac{ x }{ 100 }=>1900=25x=>\frac{ 1900 }{ 25 }=76=x=>\frac{ 76 }{ 100 }=>76percent\]

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

No problem